OpenStudy (anonymous):

Suppose that f(x)=sum_{n=0}^{infty} for all of x. If f is an odd function, show that c sub n =0 for all even numbers n.

OpenStudy (anonymous):

Suppose that \[f(x)=\sum_{n=0}^{\infty} \] for all of x. If f is an odd function, show that c sub n=0 for all even numbers n.

OpenStudy (anonymous):

so f(x) = 0?

OpenStudy (anonymous):

There is no f(x) after the summation notation. Just exactly how I typed it here. It's asking if we have the summation from n=0 to infinity of any odd function, then show that c of n = 0 for all even numbers of n.

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Is this function a constant with respect to \(x\)?

OpenStudy (anonymous):

yeah that makes no sense.

OpenStudy (anonymous):

Well at least I don't feel as bad not knowing how to even begin. :(

OpenStudy (anonymous):

Here's the copy of my worksheet....

OpenStudy (anonymous):

no the n=0 on the bottom and the infinity on top

OpenStudy (anonymous):

\[f(x)=\sum_{n=0}^{\infty}c _{n} x^n\]

OpenStudy (anonymous):

ok Then how do I show that c sub n = 0 for all even numbers n

OpenStudy (anonymous):

Proof by contradiction.

OpenStudy (anonymous):

When you plug in \(-x\), then you'll see it is different from \(-f(x)\).

OpenStudy (anonymous):

And even power of \(n\) will get rid of the negative.

OpenStudy (anonymous):

Suppose that \[f(x) = \sum_{n = 0}^{\infty}c _{n} x ^n\] for all x. Show that if f is an odd function, \[c _{0} = c _{2}=...=0\] for even \[c _{1}=c _{3}=...=0\]

OpenStudy (anonymous):

This is the question...yes, I need to prove that if it's an odd function all even numbers of n will make c sub n = 0. Can someone show me how to prove that?